# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

### v2

@@ Abstract
Efficient decoding of quantum codes is crucial for achieving high-performance quantum error correction. In this paper, we introduce the concept of approximate degenerate decoding and integrate it with ordered statistics decoding (OSD). Previously, we proposed a reliability metric that leverages both hard and soft decisions from the output of belief propagation (BP), which is particularly useful for identifying highly reliable subsets of variables. Using the approach of reliable subset reduction, we reduce the effective problem size. Additionally, we identify a degeneracy condition that allows high-order OSD to be simplified to order-0 OSD. By integrating these techniques, we present an ADOSD algorithm that significantly improves OSD efficiency in the code capacity noise model. We demonstrate the effectiveness of our BP+ADOSD approach through extensive simulations on a varity of quantum codes, including generalized hypergraph-product codes, topological codes, lift-connected surface codes, and bivariate bicycle codes. The results indicate that the BP+ADOSD decoder outperforms existing methods, achieving higher error thresholds and enhanced performance at low error rates. Additionally, we validate the efficiency of our approach in terms of computational time, demonstrating that ADOSD requires, on average, the same amount of time as two to three BP iterations on surface codes at a depolarizing error rate of around $1\%$. All the proposed algorithms are compared using single-threaded CPU implementations.
@@ I Introduction
The decoding problem for quantum codes with high degeneracy is particularly challenging for BP, as seen in topological codes [23, 24, 25, 26, 27, 28, 29]. To address these challenges, several remedies for the BP algorithm have been introduced, including techniques such as random perturbation [9], enhanced feedback [10, 11], check-matrix augmentation [12], message normalization and offset [13, 14], and trapping set analysis [30, 31]. These methods aim to improve the decoding accuracy and convergence of the BP algorithm. A recently developed modified BP4 algorithm, known as MBP4, incorporates additional memory effects and has demonstrated the ability to handle topological codes [32].
We investigate ordered statistic decoding (OSD) as a complementary approach to BP in quantum coding theory. Originally proposed by Fossorier and Lin [33] in classical coding theory for approximating maximum-likelihood decoding [34, 35, 36, 37, 38, 39, 40, 41, 42], OSD sorts reliability measures of symbols to find probable solutions while maintaining controllable complexity [33, 43, 44, 45, 46]. Syndrome-based OSD has been used when BP fails to identify a valid error [47] and is particularly effective for degenerate quantum codes [43].
The core operation of OSD is Gaussian elimination on an $m\times n$ parity-check matrix, identifying unreliable bits with a complexity of $O(nm^{2})$. Order-$w$ OSD flips up to $w$ reliable bits to generate error candidates, but this incurs high computational costs. To mitigate complexity, Fossorier et al.[47] proposed candidate pruning, and various classical techniques further optimized high-order OSD by reducing the number of candidates examined[48, 49, 50, 51, 52, 53].
[…]
However, there is a great deal of room for improvement over the current postprocessing techniques in the literature under the code capacity noise model. These methods primarily rely on binary BP to handle quaternary errors, neglecting error correlations, which suggests that binary decoders are inherently suboptimal. In contrast, the quaternary AMBP4 decoder (the adaptive version of MBP4) from [32] has been shown to outperform all the aforementioned decoders across various quantum codes, including topological codes [25], BB codes [56], and generalized hypergraph-product (GHP) codes [43, 44], without requiring postprocessing. For supporting examples, please refer to Figure 7 in Section VI.
OSD serves as a complementary strategy to BP, addressing cases where BP fails. As a result, BP-OSD is naturally expected to outperform AMBP4. In this paper, we propose a BP4-based OSD to achieve this improvement and introduce an efficient OSD algorithm for quantum codes.
First, the sorting process plays a pivotal role in OSD [33]. For efficient implementation, we consider binary representations of an $n$-qubit quantum code and thus an $n$-qubit Pauli error variable can be represented by $2n$ binary variables. Previously we proposed a reliability metric that takes into account the hard-decision history obtained during BP4 iterations and incorporates soft messages from the last iteration [1]. The $2n$ binary error variables of an error vector are then sorted according to the reliability metric. Specifically, the length of the last run in a bit’s hard-decision history is the primary term in our metric. Through simulations, we demonstrate that incorporating the hard-decision history to determine the reliability order for OSD enhances decoding performance. Our OSD algorithm based on this reliability metric will be denoted as OSD4.
To develop an efficient OSD algorithm, it is crucial to reduce the problem size in OSD and minimize the computational burden of Gaussian elimination. To address this challenge, we propose an approach called reliable subset reduction. In the code capacity error decoding problem of an $[[n,k]]$ quantum code, we need to solve a linear system of equations with $2n$ binary variables and $n-k$ independent constraints. The core idea is that if the solution for a subset of the variables can be determined with confidence, the linear system can be reduced, thereby lowering the computation burden of both Gaussian elimination and the OSD process. So the primary task is to identify a highly reliable subset. An excellent choice are the bits that exhibit nearly constant hard decision history vectors during BP4 iterations and show high soft reliability in the final iteration. At low error rates, there is a large portion of highly reliable bits. Eliminating these bits would significantly accelerate the OSD process. Moreover, resources can be allocated to the less reliable parts, similar to the approaches in [43, 44]. We will demonstrate that in our simulations of surface codes with distances from 11 to 15, the problem sizes are reduced to less than 30% of the original in over 90% of the trials at an error rate of around 2%. We note that the ambiguity clustering decoder [57] and LSD [58] operate at a different stage from our approach, and these algorithms could be applied after reliable subset reduction.
Next, we consider approximate degenerate OSD in quantum codes. A typical OSD generates a list of error candidates for a given error syndrome and selects the one with minimum weight, which is optimal for classical codes. For quantum codes, however, the optimal decoding criterion involves choosing the error coset with the highest probability, a criterion that is challenging to implement in practice. To address this, we consider approximate degenerate decoding in OSD by focusing only on relatively low-weight terms within a coset.
@@
Additionally, we introduce a degenerate error candidate pruning condition for stabilizer codes. We show that certain bit-flips in OSD correspond to multiplying stabilizers, resulting in degenerate error candidates. These bit-flips may alter the error weight but do not change the error coset, allowing them to be ignored in approximate degenerate OSD. To improve OSD efficiency, the minimum weight decoding criterion is applied. A key observation is that if all bit-flips correspond to multiplying a stabilizer, high-order OSD is unnecessary. We note that a degeneracy condition is also utilized in ambiguity clustering to facilitate subproblem resolution [57].
Integrating all the mentioned techniques, we propose an approximate degenerate OSD algorithm, abbreviated as ADOSD4. We conduct simulations of MBP4 and our OSD schemes OSD4 and ADOSD4 on several quantum codes under depolarizing errors in the code capacity noise model, including BB codes [56], rotated toric codes and rotated surface codes [25], (6.6.6) and (4.8.8) color codes [28], twisted XZZX codes on a torus [59], GHP codes [43, 44], and lift-connected surface (LCS) codes [60].
@@
The results demonstrate that our proposed schemes outperform previous approaches in the literature, both in terms of error thresholds for topological codes and low error rate performance (logical error rate around $10^{-6}$). Our MBP4+OSD4 and MBP4+ADOSD4 decoders outperform other binary BP-OSD schemes in these quantum codes. We achieve thresholds of approximately 17.5%–17.7% for the toric, surface, and twisted XZZX codes, and thresholds of 15%–15.42% for the (4.8.8) and (6.6.6) color codes, as summarized in Table III. They even surpass the minimum weight perfect matching (MWPM) decoder on surface and toric codes [69, 70].
Finally, we analyze the time efficiency of our OSD schemes for these quantum codes using single-threaded CPU implementations. The simulation results, summarized in Table IV, show that the time consumption of ADOSD4 for surface codes with distances from 11 to 15 is approximately 2% to 4% of that for order-2 OSD4 with similar decoding performance at logical error rate around $10^{-6}$. Moreover, the average time consumption per codeword for ADOSD4 is approximately equivalent to 2 to 3 BP4 iterations, whereas a successful BP trial typically takes 2 to 3 iterations on average. This suggests that ADOSD4 is an excellent complement to BP, especially since ADOSD4 is used only when BP fails.
@@
This paper is organized as follows. We introduce stabilizer codes in Section II. Section III discusses reliability metrics and the associated OSD algorithm. The reliable subset reduction method is presented in Section IV. We then cover approximate degenerate decoding in Section V, and provide simulation results in Section VI. Finally, we conclude in Sec. VII.
@@ II Quantum stabilizer codes
Every nonidentity Pauli operator in $\mathcal{G}_{n}$ has eigenvalues ±$\pm 1$. Any two Pauli operators either commute or anticommute with each other. The weight of a Pauli operator $g\in{\cal G}_{n}$, denoted $\mathrm{wt}\left(g\right)$, refers to the number of its nonidentity components.
@@
A stabilizer group $\mathcal{S}$ is an Abelian subgroup in $\mathcal{G}_{n}$ such that $-I^{\otimes n}\not\in\mathcal{S}$ [7]. Suppose that $\mathcal{S}$ is generated by $n-k$ independent generators. Then $\mathcal{S}$ defines an $[[n,k,d]]$ stabilizer code $\mathcal{C}(\mathcal{S})$ that encodes $k$ logical qubits into $n$ physical qubits:
$\mathcal{C}(\mathcal{S})=\left\{\ket{\psi}\in\mathbb{C}^{2^{n}}:g\ket{\psi}=\ket{\psi},~{}\forall\,g\in\mathcal{S}\right\}.$
@@
A syndrome decoding problem can be stated as follows: given an error syndrome of an unknown Pauli error $e$, find the most probable $\hat{e}\in{\cal G}_{n}$ that matches the syndrome. An error estimate $\hat{e}$ is considered valid if it matches the syndrome or $\hat{e}e\in{\cal S}$. Two decoding criteria are typically considered: minimum weight decoding and degenerate maximum likelihood decoding [61, 62, 63]. In minimum weight decoding, the error $\hat{e}$ with the smallest weight is chosen, while in degenerate maximum likelihood decoding, the error $\hat{e}$ whose coset $\hat{e}\mathcal{S}$ has the minimum coset probability is selected. However, degenerate maximum likelihood decoding has exponential complexity and is impractical due to the need for coset enumeration. We propose the following criterion.
@@
Note that 00-ADD reduces to minimum weight decoding, while $n$-ADD corresponds precisely to degenerate maximum likelihood decoding.
@@ II-A Binary representations
$\varphi(g)\triangleq{\mathbf{g}}=\left[\begin{array}[]{c|c}{\mathbf{g}}^{X}&{\mathbf{g}}^{Z}\end{array}\right]=\left[\begin{array}[]{c|c}{\mathbf{g}}^{X}_{1}\cdots{\mathbf{g}}^{X}_{n}&{\mathbf{g}}^{Z}_{1}\cdots{\mathbf{g}}^{Z}_{n}\end{array}\right],$
where ${\mathbf{g}}^{X}\in\{0,1\}^{n}$ and ${\mathbf{g}}^{Z}\in\{0,1\}^{n}$ are the indicator vectors of $X$ and $Z$ components of $g$, respectively, such that $\left[\begin{array}[]{c|c}{\mathbf{g}}_{j}^{X}&{\mathbf{g}}_{j}^{Z}\end{array}\right]=\varphi(M_{j}).$
A check matrix of ${\cal C}({\cal S})$ corresponding a set of stabilizer generators $\{g_{i}\}_{i=1}^{m}$ is an $m\times 2n$ binary matrix ${\mathbf{H}}$ with rows $\varphi(g_{1}),\dots,\varphi(g_{m})$.
An error $e\in{\cal G}_{n}$ can be represented by ${\mathbf{e}}=\varphi(e)$ and its error syndrome is
@@
The decoding problem is to solve the system of linear equations (2) with $2n$ binary variables in ${\mathbf{e}}\in\{0,1\}^{2n}$ that are most probable. Note that ${\mathbf{H}}$ is of rank $n-k$ so the degree of freedom of this binary system is $n+k$.
An error $e\in{\cal G}_{n}$ is classified as a logical error for a decoder if the decoder fails or outputs $\hat{e}\in{\cal G}_{n}$ such that
${\mathbf{H}}\Lambda({\mathbf{e}}+\hat{{\mathbf{e}}})^{\top}=\mathbf{0}$ (3)
@@
${\mathbf{L}}\Lambda({\mathbf{e}}+\hat{{\mathbf{e}}})^{\top}\neq\mathbf{0}.$ (4)
@@ II-B Belief Propagation (BP) Decoding
Given a check matrix ${\mathbf{H}}$, a syndrome ${\mathbf{s}}$, and a depolarizing error rate $\epsilon$, BP performs the following steps:
@@
    An initial error distribution ${\mathbf{p}}_{i}\in\mathbb{R}^{4}$ for qubit $i\in\{1,2,\dots,n\}$ is chosen for BP.
@@
    The belief distribution vector ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ will be updated by BP, using the messages of parity-check satisfaction and the channel statistics ${\mathbf{p}}_{i}$ for $i=1,2,\ldots,n$.

### v3

@@ Abstract
Efficient and scalable decoding of quantum codes is essential for high-performance quantum error correction. In this work, we introduce Reliable Subset Reduction (RSR), a reliability-driven preprocessing framework that leverages belief propagation (BP) statistics to identify and remove highly reliable qubits, substantially reducing the effective problem size. Additionally, we identify a degeneracy condition that allows high-order OSD to be simplified to order-0 OSD. By integrating these techniques, we present an ADOSD algorithm that significantly improves OSD efficiency. Our BP+RSR+ADOSD framework extends naturally to circuit-level noise and can handle large-scale codes with more than $10^{4}$ error variables. Through extensive simulations, we demonstrate improved performance over MWPM and Localized Statistics Decoding for a variety of CSS and non-CSS codes under the code-capacity noise model, and for rotated surface codes under realistic circuit-level noise. At low physical error rates, RSR reduces the effective problem size to less than 5%, enabling higher-order OSD with accelerated runtime. These results highlight the practical efficiency and broad applicability of the BP+ADOSD framework for both theoretical and realistic quantum error correction scenarios.
@@ I Introduction
The decoding problem for highly degenerate quantum codes, such as topological codes [23, 24, 25, 26, 27, 28, 29], remains particularly challenging for belief propagation (BP). A variety of remedies have been proposed to improve BP convergence and accuracy, including random perturbation [9], enhanced feedback [10, 11], check-matrix augmentation [12], message normalization and offsets [13, 14], and trapping-set analysis [30, 31]. A notable advance is MBP4 [32], which introduces memory effects and enables BP to decode many topological codes.
When BP alone is insufficient, postprocessing decoders such as ordered statistic decoding (OSD) are commonly employed. OSD was originally developed in classical coding theory as a controllable-complexity approximation to maximum-likelihood decoding [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], and later adapted to syndrome-based and quantum decoding [43, 44]. Its core operation is Gaussian elimination on an $m\times n$ parity-check matrix, followed by flipping unreliable bits, but high-order OSD incurs significant computational cost [45, 46, 47, 48, 49, 50].
In the quantum regime, several BP–OSD hybrids and variants have been proposed [51, 52, 53], along with divide-and-conquer approaches such as ambiguity clustering (AC) [54] and localized statistics decoding (LSD) [55]. These methods attempt to reduce the Gaussian elimination burden by partitioning the problem.
Despite these developments, most existing postprocessing schemes under the code-capacity noise model rely on binary BP applied to quaternary quantum errors, thereby discarding $X/Z$ correlations and leading to intrinsic suboptimality. In contrast, the quaternary AMBP4 decoder [32] directly exploits these correlations and has been shown to outperform all of the above decoders on a wide range of quantum codes, including topological codes [25], BB codes [56], and generalized hypergraph-product codes [44, 51]. However, AMBP4 still fails on certain important code families, such as lift-connected surface (LCS) codes [57], motivating the need for a more powerful postprocessing framework.
In this paper, we investigate postprocessing techniques for enhancing the performance of BP. When BP fails to converge to a valid global solution, it still provides rich statistical information about the error configuration. In particular, motivated by density evolution arguments for sparse LDPC codes [58, 59], most error variables are typically resolved with high confidence, except for a small number of unsettled variables associated with trapping sets [30]. Consequently, a large fraction of variables can be regarded as already determined and effectively part of the solution.
Based on this observation, we propose reliable subset reduction (RSR). In the code-capacity decoding problem of an $[[n,k]]$ quantum code, decoding amounts to solving a linear system with $2n$ binary variables subject to nkn-k independent constraints. The central idea of RSR is that if a subset of variables can be identified as highly reliable, the linear system can be reduced by fixing those variables, thereby substantially lowering the computational burden of subsequent postprocessing steps, such as Gaussian elimination.
To identify such a reliable subset, we exploit both the hard-decision history and the final soft information produced by BP4 [1]. Variables whose hard decisions remain nearly constant across BP iterations and whose final beliefs have high confidence are classified as reliable. At low physical error rates, a large fraction of variables satisfy these criteria, so eliminating them dramatically reduces the effective problem size and accelerates downstream postprocessing decoders.
We apply OSD after RSR, using the reliabilities produced by RSR to order the remaining unreliable variables. The resulting OSD variant, which is driven by BP4 statistics, is denoted as OSD4.
We next consider approximate degenerate OSD for quantum codes. In typical OSD, a list of error candidates consistent with a given syndrome is generated and the minimum-weight candidate is selected. For quantum codes, however, the optimal criterion is to choose the most probable error coset, which is generally intractable. To address this, we adopt an approximate degenerate decoding strategy in which only low-weight representatives within each coset are considered. When RSR is applied, the resulting reduced problem makes such approximate degenerate decoding feasible, in a manner analogous to the ambiguity clustering (AC) approach [54]. For computational efficiency, we nevertheless use minimum-weight selection among the retained candidates in our simulations.
@@
Additionally, we introduce a degeneracy-aware pruning condition for stabilizer codes. We show that certain OSD bit flips correspond to multiplying stabilizers and therefore generate degenerate error candidates: although their weights may differ, they belong to the same error coset. Such flips can be ignored in approximate degenerate OSD. A key observation is that when all candidate flips correspond to the trivial logical coset (i.e., zero logical syndrome), high-order OSD is unnecessary.
By combining these ideas, we obtain an approximate degenerate OSD algorithm, which we denote as ADOSD4.
Finally, there has been substantial recent progress in circuit-level decoding [60, 61, 62, 56, 63]. A variety of decoders have been developed for circuit-level noise models, including BP-OSD-CS, GDG, AC [54], LSD [55], and ordered Tanner forest [64]. Our BP–OSD framework can also be extended to handle both data and syndrome errors under the phenomenological noise model [65, 66], using techniques proposed in [67, 68, 69], as demonstrated in our recent work [70].
In this paper, we further show that our approach can be directly applied to the full circuit-level noise model through the detector error model (DEM) generated by STIM [71]. The resulting binary DEM decoding problem can be handled using the proposed BP–RSR–ADOSD procedure.
@@
We conduct simulations of MBP4 and our OSD schemes OSD4 and ADOSD4 on several quantum codes under depolarizing errors in the code capacity noise model, including BB codes [56], rotated toric codes and rotated surface codes [25], (6.6.6) and (4.8.8) color codes [28], twisted XZZX codes on a torus [72], GHP codes [44, 51], and LCS codes [57]. The results demonstrate that our proposed schemes outperform previous approaches in the literature, both in terms of error thresholds for topological codes and low error rate performance (logical error rate around $10^{-6}$).
In DEM circuit-level setting for rotated surface codes, we further show that MBP4+ADOSD4 consistently outperforms LSD. When both $X$- and $Z$-type detector variables are used, MBP4+ADOSD4 also surpasses the minimum-weight perfect-matching (MWPM) decoder [86, 87] and achieves an improved threshold of approximate $0.76\%$.
Beyond improved decoding performance, a central outcome of our simulations is the strong reduction effect produced by reliable subset reduction (RSR): at low physical error rates, RSR typically reduces the effective decoding problem size by one to two orders of magnitude (e.g., below $5\%$ of the original size for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$ in surface-code DEM problems). This dramatic dimensionality reduction enables higher-order OSD to be applied efficiently even for problems with more than $10^{4}$ error variables, a regime that is far beyond the reach of conventional OSD-based postprocessing. As a result, our MBP–RSR–ADOSD approach achieves strong threshold performance for topological codes and substantially improved logical error rates in the low-error regime.
@@
This paper is organized as follows. We introduce stabilizer codes in Section II. The reliable subset reduction method and the associated reliability metrics are presented in Section III. We then introduce the associated OSD algorithm in Section IV. Approximate degenerate decoding is covered in Section V. The MBP-ADOSD procedure is extended to circuit-level decoding via DEM in Section VI, and simulation results are provided in Section VII. Related work is discussed in Section VIII. Finally, we conclude in Section IX.
@@ II Quantum stabilizer codes
Every nonidentity Pauli operator in $\mathcal{G}_{n}$ has eigenvalues $\pm 1$. Any two Pauli operators either commute or anticommute with each other. The weight of a Pauli operator $g\in{\cal G}_{n}$, denoted $\mathrm{wt}\left(g\right)$, refers to the number of its nonidentity components.
@@
A stabilizer group $\mathcal{S}$ is an Abelian subgroup in $\mathcal{G}_{n}$ such that $I^{\otimes n}\not\in\mathcal{S}$ [7]. Suppose that $\mathcal{S}$ is generated by nkn-k independent generators. Then $\mathcal{S}$ defines an $[[n,k,d]]$ stabilizer code $\mathcal{C}(\mathcal{S})$ that encodes $k$ logical qubits into $n$ physical qubits:
[…]
@@
A syndrome decoding problem is defined as follows: given a measured syndrome corresponding to an unknown Pauli error $e$, a decoder outputs an estimate $\hat{e}\in\mathcal{G}_{n}$. An error estimate $\hat{e}$ is considered valid if it reproduces the measured syndrome. The decoder is said to succeed if the estimate is degenerate with the true error, i.e.,
$\hat{e}e\in{\cal S},$
or equivalently, $\hat{e}e$ commutes with all elements of the normalizer group $N({\cal S})$. Otherwise, the decoder produces a logical error.
Two decoding criteria are typically considered: minimum weight decoding and degenerate maximum likelihood decoding [74, 75, 76]. In minimum weight decoding, the error $\hat{e}$ with the smallest weight is chosen, while in degenerate maximum likelihood decoding, the error $\hat{e}$ whose coset $\hat{e}\mathcal{S}$ has the minimum coset probability is selected. However, degenerate maximum likelihood decoding has exponential complexity and is impractical due to the need for coset enumeration. We propose the following criterion.
@@
Note that 0-ADD reduces to minimum weight decoding, while $n$-ADD corresponds precisely to degenerate maximum likelihood decoding.
@@ II-A Binary representations
$\varphi(g)\triangleq{\mathbf{g}}=\left[\begin{array}[]{c|c}{\mathbf{g}}^{X}&{\mathbf{g}}^{Z}\end{array}\right]=\left[\begin{array}[]{c|c}a_{1}\cdots a_{n}&b_{1}\cdots b_{n}\end{array}\right]\in\{0,1\}^{2n},$
where ${\mathbf{g}}^{X}=(a_{1},\dots,a_{n})\in\{0,1\}^{n}$ and ${\mathbf{g}}^{Z}=(b_{1},\dots,b_{n})\in\{0,1\}^{n}$ are the indicator vectors of $X$ and $Z$ components of $g$, respectively, such that $\varphi(M_{j})=\left[\begin{array}[]{c|c}a_{j}&b_{j}\end{array}\right].$
A check matrix of ${\cal C}({\cal S})$ corresponding to a set of stabilizer generators $\{g_{i}\}_{i=1}^{m}$ is an $m\times 2n$ binary matrix $\mathbf{H}$} with rows $\varphi(g_{1}),\dots,\varphi(g_{m})$.
An error $e\in{\cal G}_{n}$ can be represented by ${\mathbf{e}}=\varphi(e)\in\{0,1\}^{2n}$ and its error syndrome is
@@
The decoding problem is to solve the system of linear equations (2) with $2n$ binary variables in ${\mathbf{e}}\in\{0,1\}^{2n}$ that are most probable. Note that $\mathbf{H}$} is of rank nkn-k so the degree of freedom of this binary system is $n+k$.
An error $e\in{\cal G}_{n}$ is classified as a logical error for a decoder if the decoder fails to produce a valid estimate, or outputs an estimate $\hat{e}\in{\cal G}_{n}$ such that
${\mathbf{H}}\Lambda{\mathbf{e}}^{\top}={\mathbf{H}}\Lambda\hat{{\mathbf{e}}}^{\top}$ (3)
@@
${\mathbf{L}}\Lambda{\mathbf{e}}^{\top}\neq{\mathbf{L}}\Lambda\hat{{\mathbf{e}}}^{\top}.$ (4)
We define the logical syndrome of an error $\mathbf{e}$} as
${\mathbf{L}}\Lambda{\mathbf{e}}^{\top}\in\{0,1\}^{2k\times 1}.$ (5)
@@ II-B Belief Propagation (BP) Decoding
Given a check matrix $\mathbf{H}$}, a syndrome $\mathbf{s}$}, and the prior error distributions ${\mathbf{p}}_{i}\in\mathbb{R}^{4}$ for qubit $i\in\{1,2,\dots,n\}$, BP performs the following steps:
@@
    The belief distribution vector ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ is updated by BP, using the parity-check messages and the prior ${\mathbf{p}}_{i}$, for $i=1,2,\ldots,n$.
@@
[…]
These two steps are iterated for at most $\mathrm{T}$} iterations, where $\mathrm{T}$} is chosen in advance. If no valid estimate is obtained after $\mathrm{T}$} iterations, BP declares a decoding failure.
III The Reliable Subset Reduction Method
In practice, a BP decoder typically either converges within a small number of iterations, on the order of $O(\log\log n)$ on average, or remains unsettled until the maximum number of iterations is reached. In the latter case, although BP fails to converge globally, most error variables have already been decided with high confidence. From the BP output, a large fraction of variables are therefore very reliable and can be regarded as part of the solution. The remaining unsettled variables are usually associated with trapping sets [30], and additional post-processing techniques are required to resolve them.
In this section, we estimate the expected size of these unsettled error variables using density evolution [58, 59].
III-A Density-Evolution-Guided Reliability Heuristic
We consider a family of regular LDPC codes over a quaternary symmetric channel and use the syndrome-based Gallager-A algorithm [15] as a heuristic approximation of quantum stabilizer decoding under depolarizing noise. Density evolution is then employed to characterize the behavior of this iterative decoder.
Gallager-A is a hard-decision message-passing algorithm defined on a Tanner graph. In our setting, variable nodes $v_{j}$ represent Pauli errors on physical qubits, while check nodes $c_{i}$ represent stabilizer generators. Since a Pauli operator either commutes or anticommutes with a stabilizer, all messages are binary: $1$ indicates anti-commutation and 0 indicates commutation.
We consider a syndrome-based quaternary–binary variant of Gallager-A suitable for CSS-type stabilizer codes, where each stabilizer contains only $X$ or $Z$ operators. Given a check matrix $\mathbf{H}$} and syndrome ${\mathbf{s}}=(s_{1},\dots,s_{m})$, the decoding proceeds as follows.
  1. 1.
    Initialization: Set estimated error $\hat{E}_{j}=I$ for all qubits and initialize all variable-to-check messages
    $m_{v_{j}\to c_{i}}^{(0)}=0.$
  2. 2.
    Check update: At iteration $\ell$, each check node computes a recommendation message
    $m_{c_{i}\to v_{j}}^{(\ell)}=s_{i}\oplus\bigoplus_{k\in{\cal I}(c_{i})\setminus\{j\}}m_{v_{k}\to c_{i}}^{(\ell)}.$